# Movement-Prediction-Adjusted Naïve Forecast

### v8

@@ 3.2 Results and Analysis
Because the directional accuracy for the out-of-sample set is predefined in each simulation process, the value of $\theta_{\text{out-of-sample}}^{*}$ can be derived directly via Eq. (14), and only the value of $\bar{\epsilon}_{\text{out-of-sample}}$ needs to be estimated. Figure 5 illustrates the distribution of the values in the $\boldsymbol{\bar{\epsilon}}_{\text{set}}$, generated based on the sliding windows across the first-order differences of the in-sample set. The vertical reference line in Figure 5 represents the actual value of $\bar{\epsilon}_{\text{out-of-sample}}$. In the four cases, the minimum value of $\boldsymbol{\bar{\epsilon}}_{\text{set}}$ is generally an underestimate of $\bar{\epsilon}_{\text{out-of-sample}}$, except for the second case, where the synthetic time series is a random walk with a decreasing trend of linear variance. In general, the estimated value remains close to the true value. Under the assumption that the directional accuracy is consistent across the in-sample and out-of-sample sets, this estimation ensures that inequality (16) is satisfied.
Figure 5: Comparison of the true $\bar{\epsilon}_{\text{out-of-sample}}$ with $\boldsymbol{\bar{\epsilon}}_{\text{set}}$: Type 1, constant variance ($\sigma_{0}^{2}=1$); Type 2, linear variance trend ($k=4.95$); Type 3, cyclic variance ($a=7.77$); and Type 4, randomly varying variance ($\xi^{2}=920$).
Furthermore, comparisons of the two types of naïve forecasts based on the four evaluation metrics at low accuracy levels are illustrated in Figures 6, 7, 8, and 9. Each box plot in each figure represents the distribution of 100 independent values for one metric at a specific directional accuracy level, with the metric value of the original naïve forecast shown as a reference line. According to Figure 6, at low accuracy levels, such as 0.51, the RMSE of the adjusted naïve forecast is comparable to or slightly better than that of the original naïve forecast, as shown by the proximity of the median of each box plot to the reference line. As the accuracy increased beyond 0.52, most of the RMSE values fell below the baseline.
Figure 6: RMSEs of the adjusted naïve forecast across different directional accuracy levels (0.50–0.56) for four synthetic time series: Type 1, constant variance; Type 2, linear variance trend; Type 3, cyclic variance; and Type 4, randomly varying variance.
Figure 7: MAEs of the adjusted naïve forecast across different directional accuracy levels (0.50–0.56) for four synthetic time series: Type 1, constant variance; Type 2, linear variance trend; Type 3, cyclic variance; and Type 4, randomly varying variance.
Figure 8: MAPEs of the adjusted naïve forecast across different directional accuracy levels (0.50–0.56) for four synthetic time series: Type 1, constant variance; Type 2, linear variance trend; Type 3, cyclic variance; and Type 4, randomly varying variance.
Figure 9: sMAPEs of the adjusted naïve forecast across different directional accuracy levels (0.50–0.56) for four synthetic time series: Type 1, constant variance; Type 2, linear variance trend; Type 3, cyclic variance; and Type 4, randomly varying variance.
Because the adjusted naïve forecast consistently outperforms the original naïve forecast at low accuracy levels in terms of the MAE, MAPE, and sMAPE, the Wilcoxon signed rank test was conducted to specifically evaluate whether the median differences between the RMSE values of the adjusted naïve forecast and those of the original naïve forecast significantly deviated from zero for accuracy levels ranging from 0.50–0.56 (Li and Johnson,, 2014). As summarized in Table 2, at a directional accuracy level of 0.51, the adjusted naïve forecast shows mixed results with respect to statistical significance. Although significant improvements were observed over the original naïve forecast for the Type 4 series (randomly varying variance, p = 0.002), the results for the Type 1 (constant variance, p = 0.915), Type 2 (linear variance trend, p = 0.781), and Type 3 (cyclic variance, p = 0.812) series were not statistically significant, indicating that the adjusted naïve forecast under an accuracy level of 0.51 may not consistently produce meaningful improvements across all types of time series. As the accuracy increased beyond 0.51, significant reductions in the RMSE (p < 0.001) were observed for all types of time series, thereby demonstrating the reliability and effectiveness of the adjusted naïve forecast at higher accuracy levels.
Table 2: RMSE results of the adjusted naïve forecast across directional accuracy levels (0.50–0.56) for four synthetic time series: Type 1, constant variance; Type 2, linear variance trend; Type 3, cyclic variance; and Type 4, randomly varying variance.
Prediction Type 1 Type 2 Type 3 Type 4
RMSE Range p-value RMSE Range p-value RMSE Range p-value RMSE Range p-value
Original Naïve 0.961 - 0.00762 - 1.355 - 21.034 -
Acc = 0.50 0.961 - 0.00762 - 1.355 - 21.034 -
Acc = 0.51 0.960 $\pm$ 0.0004 0.915 0.00762 $\pm$ 0.000003 0.781 1.355 $\pm$ 0.0005 0.812 21.032 $\pm$ 0.0053 0.002
Acc = 0.52 0.960 $\pm$ 0.0008 <0.001 0.00762 $\pm$ 0.000007 <0.001 1.355 $\pm$ 0.0008 <0.001 21.024 $\pm$ 0.0102 <0.001
Acc = 0.53 0.959 $\pm$ 0.0012 <0.001 0.00761 $\pm$ 0.000010 <0.001 1.354 $\pm$ 0.0015 <0.001 21.015 $\pm$ 0.0144 <0.001
Acc = 0.54 0.959 $\pm$ 0.0017 <0.001 0.00761 $\pm$ 0.000014 <0.001 1.353 $\pm$ 0.0020 <0.001 20.995 $\pm$ 0.0199 <0.001
Acc = 0.55 0.957 $\pm$ 0.0020 <0.001 0.00760 $\pm$ 0.000020 <0.001 1.352 $\pm$ 0.0028 <0.001 20.979 $\pm$ 0.0244 <0.001
Acc = 0.56 0.956 $\pm$ 0.0023 <0.001 0.00758 $\pm$ 0.000021 <0.001 1.351 $\pm$ 0.0032 <0.001 20.957 $\pm$ 0.0278 <0.001
According to Figures 10, 11, 12, and 13, at higher accuracy levels ranging from 0.55–1.00, the adjusted naïve forecast consistently outperforms the original naïve forecast in terms of the four metrics. In particular, the improvement over the original naïve forecast becomes substantial as the directional accuracy improves. This finding demonstrates that the proposed method can effectively translate highly accurate movement predictions into improved point forecasts in the context of forecasting random walk-like time series. However, the improvement over the original naïve forecast has a limit. The proposed method allows a random walk-like series to be predicted to a certain extent, but the predicted value cannot be infinitely close to the actual value.
Figure 10: RMSEs of the adjusted naïve forecast across different directional accuracy levels (0.55–1.00) for four synthetic time series: Type 1, constant variance; Type 2, linear variance trend; Type 3, cyclic variance; and Type 4, randomly varying variance.
Figure 11: MAEs of the adjusted naïve forecast across different directional accuracy levels (0.55–1.00) for four synthetic time series: Type 1, constant variance; Type 2, linear variance trend; Type 3, cyclic variance; and Type 4, randomly varying variance.
Figure 12: MAPEs of the adjusted naïve forecast across different directional accuracy levels (0.55–1.00) for four synthetic time series: Type 1, constant variance; Type 2, linear variance trend; Type 3, cyclic variance; and Type 4, randomly varying variance.
Figure 13: sMAPEs of the adjusted naïve forecast across different directional accuracy levels (0.55–1.00) for four synthetic time series: Type 1, constant variance; Type 2, linear variance trend; Type 3, cyclic variance; and Type 4, randomly varying variance.
In general, when the directional accuracy was 0.5, the adjusted naïve forecast did not provide an improvement over the original naïve forecast. As the directional accuracy increases, the performance of the adjusted naïve forecast improves. The simulation results also demonstrate that the effectiveness of the proposed method is unaffected by the variance type of the synthetic time series, as evidenced by the consistent performance of the adjusted naïve forecast across all synthetic series. This result is consistent with the theoretical analysis of the proposed method, which highlights its robustness and potential applicability to a wide range of time series that exhibit symmetric random walk characteristics, including those with homoscedasticity as well as high-volatility scenarios.
@@ 4 Discussion
This section discusses the implications of the findings and potential limitations of this study. First, this study demonstrates an indirect regression approach that leverages movement predictions in conjunction with naïve forecasts for time series exhibiting random walk-like behavior. The theoretical analysis and simulation results suggest that accurate movement prediction can be systematically translated into improved point forecasts. This highlights the potential value of incorporating directional information into simple baseline models and suggests that enhancing movement prediction may, in some contexts, be more effective than developing increasingly complex direct regression models.
@@
Second, in many real-world forecasting scenarios, decision-making often involves integrating predictions with heterogeneous data types, such as combining binary classifications and continuous regressions. However, such integration can be technically challenging. The proposed method provides a practical means of converting binary movement predictions into continuous point forecasts for random walk-like time series, thereby enabling more informed decision-making in applied forecasting contexts.
Third, the goal of this study is to examine the theoretical possibility of improving forecasts for random walk-like time series by incorporating directional signals rather than disputing the broader validity of the naïve forecast as the optimal benchmark. While the simulation implies that even weak directional signals can be exploited to enhance point forecasts, these improvements can be modest and context dependent. The adjusted naïve forecast serves as a second-stage method only after movement prediction is provided. Moreover, such improvements do not necessarily translate into economic gains, particularly when real-world constraints such as transaction costs, slippage, and market frictions are considered.
Finally, because the focus of this study is on the theoretical impact of directional information rather than on modeling movement predictions themselves, we rely on controlled simulations rather than empirical data. The effectiveness of the proposed adjustment depends on the consistency of the directional accuracy across different segments of the series. In practice, movement prediction models susceptible to overfitting may suffer from deteriorated performance on out-of-sample data, thereby limiting the actual benefit of incorporating such predictions into the original naïve forecast.
5 Conclusion
This study introduces a movement-prediction-adjusted naïve forecast, which is a linear combination of movement prediction and the original naïve forecast, in the challenging context of forecasting time series that exhibit symmetric random walk characteristics. The weight of the movement term is determined by two parameters derived from the in-sample data: one based on the directional accuracy of the movement prediction and the other on the mean absolute increment of the target series. The theoretical analysis of the proposed method explains how accurate movement prediction can be exploited to enhance point forecasting for a random walk-like series. It also defines a boundary criterion stating that the adjusted naïve forecast outperforms the original naïve forecast when the product of the actual optimal parameter values exceeds half the product of their estimated values. Furthermore, the simulation results demonstrate that the adjusted naïve forecast can achieve statistically significant improvements over the original naïve forecast, even with directional accuracies slightly greater than 0.50. These findings highlight the significant potential of using movement predictions to address the inherent challenges of point forecasting in a random walk-like time series.
There are several avenues for further research on the foundation of this study. First, while this research focuses on time series with symmetric random walk characteristics, future work could explore applying this movement prediction-based method to time series with nonsymmetric random walk properties, potentially broadening its applicability. Second, it is worth exploring whether the proposed forecasting method can be generalized beyond the one-dimensional case, as many real-world variables exhibit random walk behavior in a high-dimensional space. Moreover, the proposed method uses a fixed increment for simplicity. However, adapting this increment based on the availability of new data may provide more dynamic and accurate forecasts. Investigating how the increments can be adapted and whether such adaptive increments can further refine the forecasting accuracy represents a promising direction for future research.
@@ Acknowledgments
Appendix A. Algebraic Derivation of $\Delta MSE_{\text{in-sample}}$
According to Eq. (2), which presents the actual value of the target variable, the MSE expression for the original naïve forecast can be expressed as:
$MSE_{\text{in-sample}}^{\text{original-na\"{\i}ve}}$ $=\frac{1}{N}\left(\sum_{t\in\text{in-sample}}\left(y_{t}-y_{t-1}\right)^{2}\right)$ (A1)
@@ Appendix A. Algebraic Derivation of $\Delta MSE_{\text{in-sample}}$
Similarly, by substituting Eq. (2) for actual values and Eq. (4) for predicted values, the MSE of the adjusted naïve forecast can be expressed as follows:
$MSE_{\text{in-sample}}^{\text{adjusted-na\"{\i}ve}}$ $=\frac{1}{N}\left(\sum_{t\in\text{in-sample}}\left(y_{t}-\hat{y}_{t}\right)^{2}\right)$ (A2)
@@
Hence, the MSE difference between the adjusted naïve forecast and the original naïve forecast on the in-sample set can be derived via Eq. (A1) and Eq. (A2):
$\Delta\text{MSE}_{\text{in-sample}}$ $=\text{MSE}_{\text{in-sample}}^{\text{original-na\"{\i}ve}}-\text{MSE}_{\text{in-sample}}^{\text{adjusted-na\"{\i}ve}}$ (A3)
@@
Based on the law of large numbers (LLN), when both $n_{\text{in-sample}}^{\text{correct}}$ and $n_{\text{in-sample}}^{\text{incorrect}}$ are sufficiently large, the following approximations hold:
@@
$\Delta MSE_{\text{in-sample}}$ $=MSE_{\text{in-sample}}^{\text{original-na\"{\i}ve}}-MSE_{\text{in-sample}}^{\text{adjusted-na\"{\i}ve}}$ (A5)
@@
Appendix B. Algebraic Derivation of $\Delta MSE_{\text{out-of-sample}}$
When using $\theta_{\text{in-sample}}^{*}$ and $\bar{\epsilon}_{\text{in-sample}}$ as parameters for the adjusted naïve forecast on the out-of-sample set, the difference in MSE between the adjusted and original naïve forecasts on the out-of-sample set can be expressed as follows:
$\Delta\text{MSE}_{\text{out-of-sample}}$ $=\text{MSE}_{\text{out-of-sample}}^{\text{original-na\"{\i}ve}}-\text{MSE}_{\text{out-of-sample}}^{\text{adjusted-na\"{\i}ve}}$ (B1)

### v9

@@ 3.2 Results and Analysis
The out-of-sample forecasting results for the eight financial time series are illustrated in Figure 4. Given that each out-of-sample set consists of 1250 time steps and that predictions from different methods visually overlap to a large extent, only the first five predicted points from each set are displayed for clarity. Small differences between the MPANF and the other forecasting methods are observed because the overall in-sample movement prediction accuracy for each time series is just slightly above 0.55, leading to a small coefficient $\theta$ (approximately 0.1) and a modest adjustment of the naive forecast at each time step. The naive forecast, naive forecast with drift, and IMA(1,1) forecasts yield similar predictions, reinforcing their suitability as baseline methods for time series forecasting. Moreover, because the linear regression model closely mirrors the structure of the MPANF, its performance is similarly dependent on movement prediction accuracy and therefore remains close to that of the MPANF.
Figure 4: Out-of-sample forecasts by different methods.
Tables 2, 3, 4, and 5 present performance comparisons between the MPANF and the baseline methods for eight out-of-sample sets, reported in terms of the RMSE, MAE, MAPE, and sMAPE, respectively. Notably, the MPANF consistently outperforms the naive forecast, with the only exceptions being the MAPE and sMAPE for the GOOGL dataset. Specifically, the MPANF achieves the lowest RMSE in 4 out of 8 cases, the lowest MAE in 6 out of 8 cases, the lowest MAPE in 6 out of 8 cases, and the lowest sMAPE in 5 out of 8 cases, demonstrating its overall superiority over the baseline methods.
Table 2: RMSE comparison across forecasting methods for selected stocks
Stock Naive Naive + Drift IMA(1,1) LR MPANF
AAPL 2.5937 2.5925 2.5979 2.5942 2.5931
BA 5.9080 5.9187 5.9166 5.8932 5.8829
GOOGL 2.2093 2.2085 2.2108 2.2114 2.2088
GS 5.3671 5.3660 5.3727 5.4039 5.3485
HAL 0.7334 0.7341 0.7332 0.7261 0.7271
OIL 1.8909 1.8912 1.8976 1.8940 1.8864
SLB 0.9401 0.9410 0.9401 0.9331 0.9343
TSLA 8.1524 8.1525 8.1677 8.3351 8.1515
Table 3: MAE comparison across forecasting methods for selected stocks
Stock Naive Naive + Drift IMA(1,1) LR MPANF
AAPL 1.8950 1.8935 1.8975 1.8958 1.8943
BA 3.9691 3.9821 3.9750 3.9657 3.9533
GOOGL 1.5680 1.5660 1.5727 1.5746 1.5674
GS 3.9743 3.9737 3.9811 4.0021 3.9534
HAL 0.5423 0.5429 0.5439 0.5376 0.5368
OIL 1.2959 1.2981 1.3000 1.3073 1.2903
SLB 0.6901 0.6910 0.6907 0.6852 0.6843
TSLA 5.5732 5.5733 5.5775 5.8155 5.5722
Table 4: MAPE (%) comparison across forecasting methods for selected stocks
Stock Naive Naive + Drift IMA(1,1) LR MPANF
AAPL 1.4081 1.4069 1.4065 1.4079 1.4074
BA 2.0601 2.0675 2.0633 2.0585 2.0514
GOOGL 1.4357 1.4340 1.4380 1.4406 1.4359
GS 1.4037 1.4036 1.4046 1.4074 1.3967
HAL 2.3517 2.3532 2.3667 2.3277 2.3216
OIL 1.8379 1.8409 1.8430 1.8456 1.8296
SLB 2.1522 2.1532 2.1574 2.1355 2.1291
TSLA 2.9511 2.9513 2.9549 3.0534 2.9509
Table 5: sMAPE (%) comparison across forecasting methods for selected stocks
Stock Naive Naive + Drift IMA(1,1) LR MPANF
AAPL 1.4088 1.4072 1.4072 1.4088 1.4081
BA 2.0542 2.0591 2.0591 2.0500 2.0457
GOOGL 1.4352 1.4331 1.4375 1.4410 1.4354
GS 1.4035 1.4032 1.4045 1.4090 1.3966
HAL 2.3377 2.3427 2.3537 2.3051 2.3084
OIL 1.8290 1.8328 1.8336 1.8377 1.8210
SLB 2.1465 2.1503 2.1521 2.1260 2.1240
TSLA 2.9518 2.9520 2.9554 3.0767 2.9515
In addition, baseline methods such as naive with drift, IMA(1,1), or linear regression occasionally yield lower RMSE values (as observed for AAPL, GOOGL, HAL, and OIL), lower MAE values (AAPL, GOOGL), lower MAPE values (AAPL, GOOGL), and lower sMAPE values (AAPL, GOOGL, and HAL). However, the performances of these methods are inconsistent; they frequently fall short of the naive forecast in other cases.
Notably, although linear regression also uses movement prediction as an input feature and may occasionally outperform the naive forecast, it lacks a theoretical explanation for such improvement. Moreover, its performance generally lags behind that of the MPANF. This performance gap arises because linear regression is more susceptible to overfitting. While it closely fits the in-sample data, its out-of-sample performance can deteriorate significantly if the movement prediction accuracy decreases. In contrast, the coefficient determination in the MPANF is guided by theoretically derived optimal conditions rather than by data-driven minimization of in-sample errors. This design inherently introduces regularization, effectively mitigating overfitting and enhancing the method’s generalization ability.
Additionally, Figures 5, 6, 7, and 8 present the performance of each forecasting method across the four evaluation metrics. Overall, the MPANF generates more accurate and robust forecasts than the baseline methods do and consistently improves upon the naive forecast across diverse financial forecasting scenarios. These empirical results are consistent with the theoretical analysis, underscoring the method’s broad applicability to financial time series forecasting, including series characterized by heteroscedasticity and high volatility.
Figure 5: Comparison of the RMSEs of different methods.
Figure 6: Comparison of the MAEs of different methods.
Figure 7: Comparison of the MAPEs of different methods.
Figure 8: Comparison of the sMAPEs of different methods.
@@ 4 Discussion
This section discusses the implications of the findings and the potential limitations of the study.
@@
First, the results demonstrate the effectiveness of a second-stage forecasting method that combines movement predictions with naive forecasts for financial time series forecasting. Both the theoretical analysis and the empirical evidence indicate that accurate movement prediction can be systematically translated into improved point forecasts. This underscores the value of incorporating directional information into a naive forecast and suggests that, in some contexts, the indirect forecasting approach may be more effective than developing complex regression models.
@@
Second, in many real-world forecasting applications, decision-making requires integrating predictions from heterogeneous data types, such as binary classifications and continuous regressions. Such integration can be technically challenging. The proposed method offers a practical solution by converting binary movement predictions into continuous point forecasts, thereby facilitating more informed decisions in forecasting scenarios.
Third, while the empirical results show that even relatively weak directional signals can enhance point forecasts, the magnitude of such improvements is often modest and context dependent. Moreover, these gains do not necessarily translate into economic benefits, especially when real-world factors such as transaction costs, slippage, and market frictions are taken into account.
Finally, although the MPANF imposes no restrictions on the choice of movement prediction models, its effectiveness depends critically on the consistency and reliability of directional accuracy across different segments of the dataset. In practice, predicting the movements of financial time series remains challenging, and models prone to overfitting may suffer substantial degradation in out-of-sample performance. Such deterioration can reduce, or even eliminate, the practical advantages of incorporating movement predictions into the naive forecast. In other words, the applicability of the MPANF is conditional on the availability of a meaningful and robust movement prediction for the target variable.
5 Conclusions
This study introduces the MPANF, a forecast combination method that adjusts the naive forecast using the movement prediction of the target variable. The approach is particularly effective for time series in which surpassing the naive forecast is typically difficult but accurate movement prediction is available. The weight assigned to the movement term is calculated as the product of a coefficient derived from the in-sample movement prediction accuracy and the in-sample mean absolute increment. Theoretical analysis explains why this combination can yield point forecasts that are superior to the naive forecast. Empirical evaluation on eight financial time series confirms the MPANF’s ability to achieve robust improvements over the naive forecast, even when the directional accuracy is only modestly above 0.55. Together, these findings underscore the potential of leveraging movement predictions to enhance point forecasts in challenging forecasting scenarios.
Several avenues for future research emerge from this work. First, because the present evaluation was conducted under modest directional accuracy, developing more advanced movement prediction models and assessing MPANF performance at higher accuracy levels could enhance its practical value in real-world trading. Second, applying MPANF to other time series domains—particularly those where the naive forecast is hard to beat yet movement prediction is feasible—could broaden its applicability. Third, extending the MPANF beyond the univariate case would address the many forecasting tasks that arise in high-dimensional settings. Finally, while this study employs a fixed adjustment for simplicity, exploring adaptive adjustment schemes may further improve the forecasting accuracy.
@@ Acknowledgments
Appendix A. Algebraic Derivation of $\Delta MSE_{\textnormal{in-sample}}$
According to Eq. (2), the in-sample MSE expression for the naive forecast, denoted as $MSE_{\text{in-sample}}^{\text{Naive}}$, can be expressed as:
$MSE_{\text{in-sample}}^{\text{Naive}}$ $=\frac{1}{N}\left(\sum_{t\in\text{in-sample}}\left(y_{t}-y_{t-1}\right)^{2}\right)$ (A1)
@@ Appendix A. Algebraic Derivation of $\Delta MSE_{\textnormal{in-sample}}$
By substituting Eq. (2) into $y_{t}$ and Eq. (4) into $y_{t-1}$, the in-sample MSE of the MPANF, denoted as $MSE_{\text{in-sample}}^{\text{MPANF}}$, can be expressed as follows:
$MSE_{\text{in-sample}}^{\text{MPANF}}$ $=\frac{1}{N}\left(\sum_{t\in\text{in-sample}}\left(y_{t}-\hat{y}_{t}\right)^{2}\right)$ (A2)
@@
Hence, the in-sample MSE difference between the naive forecast and the MPANF, denoted as $\Delta MSE_{\text{in-sample}}$, can be derived via Eq. (A1) and Eq. (A2):
$\Delta MSE_{\text{in-sample}}$ $=MSE_{\text{in-sample}}^{\text{Naive}}-MSE_{\text{in-sample}}^{\text{MPANF}}$ (A3)
@@
Based on the law of large numbers (LLN), under idealized conditions where both $n_{\text{in-sample}}^{\text{correct}}$ and $n_{\text{in-sample}}^{\text{incorrect}}$ are sufficiently large, and the two sets of increments corresponding to correct and incorrect movement predictions are unbiased and representative of the overall increment set, the following approximations hold:
@@
$\Delta MSE_{\text{in-sample}}$ $=MSE_{\text{in-sample}}^{\text{Naive}}-MSE_{\text{in-sample}}^{\text{MPANF}}$ (A5)
@@
Appendix B. Algebraic Derivation of $\Delta MSE_{\textnormal{out-of-sample}}$
When $\theta_{\text{in-sample}}^{*}$ and $\bar{\epsilon}_{\text{in-sample}}$ are used by the MPANF on the out-of-sample set, the out-of-sample MSE difference between the naive forecast and the MPANF, denoted as $\Delta MSE_{\text{out-of-sample}}$, can be expressed as follows:
$\Delta MSE_{\text{out-of-sample}}$ $=MSE_{\text{out-of-sample}}^{\text{Naive}}-MSE_{\text{out-of-sample}}^{\text{MPANF}}$ (B1)